# Networked Information Aggregation via Machine Learning

### v1

@@ 1.1 Our Results and Techniques
We show that a predictor $f$ that is multiaccurate with respect to a collection of functions $G$ and also self-orthogonal has mean squared error that is at most that of the optimal function in $G$ [^1] For binary classification problems, it is known that multiaccuracy together with calibration is sufficient to yield accuracy competitive with the best model in $G$ Gopalan et al. (2023); Casacuberta et al. (2025). Calibration is a substantially stronger condition than self-orthogonality, and implies it. To obtain calibrated multiaccuracy we would need the learners to step outside of the class of linear models — but linear regression automatically guarantees self-orthogonality. — this is a consequence of the simple MSE decomposition that we give in Lemma 3.3:
@@
Now recall that one of the inputs to agent $i$’s model $f_{i}$ is the prediction $f_{j}$ of any agent $j$ that is a parent of $i$ in the graph. Hence, it follows from $f_{i}$’s multiaccuracy guarantee that $\mathbb{E}[f_{j}(f_{i}-y)]=0$. We show in Lemma 3.8 that this condition implies that:
@@ 1.2 Related Work
Banerjee (1992) and Bikhchandani et al. (1992) study simple models in which Bayesian agents act sequentially, each taking an action that maximizes their utility in expectation over their posterior beliefs of a single, static state. In these models, agents observe the actions of all previous agents, and the main phenomenon of interest is “herding” — i.e. when an infinite sequence of agents takes the wrong action, as the cumulative decisions of their predecessors overwhelm the strength of their own signal. Acemoglu et al. (2011) extend this model to general networks in which agents do not observe the actions of all predecessors, but only those adjacent to them in the network. They study conditions (expansion properties) under which information aggregation occurs in the infinite limit. Mossel et al. (2015, 2020) study a variant of this style of model in which agents act repeatedly rather than sequentially, and study information aggregation in equilibrium. Jadbabaie et al. (2012); Shahrampour and Jadbabaie (2013) study a related model in which agents make Bayesian updates with respect to their own signals, but then broadcast their posterior and simply average their posterior beliefs with those of their neighbors, in the style of a DeGroot (1974) model. They similarly give conditions sufficient for information aggregation.
@@
Technically, our analysis is most closely related to a recent line of work that uses conditional calibration conditions for ensembling and information aggregation Roth et al. (2023); Globus-Harris et al. (2023, 2024); Collina et al. (2025a, b). Two recent works of Collina et al. (2025b, a) in particular show how to recover the same quantitative results in the style of Aumann’s agreement theorem (Aumann, 1976; Aaronson, 2005) and its information aggregation properties Frongillo et al. (2023) without assuming full Bayesian rationality, but instead assuming only simple calibration conditions that can be tractably enforced even in sequential adversarial settings (Noarov et al., 2023; Arunachaleswaran et al., 2025). Our work is distinguished from this literature in two ways: first, this literature considers agents who engage in an extended interaction in which each agent participates multiple times, solving many learning problems across this interaction. In contrast, in our model, each agent learns only once, and then fixes their model: information aggregation needs to occur as predictions percolate through a network of agents, no pair of which might observe the same set of features. Second, this literature assumes stronger conditional calibration conditions (“multi-calibration” (Hébert-Johnson et al., 2018)) that require learning more complicated non-linear models, even to compete with linear benchmark classes. In contrast, our analysis requires only much weaker orthogonality conditions (also known as “multi-accuracy” (Hébert-Johnson et al., 2018; Kim et al., 2019)) which not only are satisfiable by linear models but are already satisfied by least squares regression, and so need no further algorithmic intervention. Our model could be used to simulate the repeated interaction studied in Collina et al. (2025a, b) by constructing a path graph in which each agent was repeatedly represented by alternating nodes — for squared error minimization over additively separable classes in the batch setting, our results would recover some of those of Collina et al. (2025a) (note however that Collina et al. (2025a) also studies more complex settings with abstract action spaces and utility functions, the online adversarial setting, and a more general weak learning condition).
@@ 2 Preliminaries
Let $[d]=\{1,2,\ldots,d\}$ be the set of indices for $d$ total features. Each agent $A_{i}\in\mathcal{A}$ is associated with a specific subset of these features, $S_{i}\subseteq[d]$. We assume data points $(x,y)$ are drawn from an underlying distribution $\mathcal{D}$, where $x\in\mathbb{R}^{d}$ is a feature vector and $y\in\mathbb{R}$ is the target label. For any agent $A_{i}$, its local view of the features is $x_{S_{i}}$, which is the sub-vector of $x$ corresponding to the features indexed by $S_{i}$.
@@
This means that the predictor’s errors $f(x)-y$ have low correlation with every function $g$ in the collection $\mathcal{G}$. If $\epsilon=0$ we simply say that $f$ is multiaccurate with respect to $\mathcal{G}$.
@@ 3.1 Structural Properties of Squared Error Regression
Next we derive a useful identity that we will make use of several times: for any pair of predictors $f$ and $g$, we can express the MSE of $f$ in terms of:
@@
This decomposition does not depend on linearity:
@@
We start by expanding $\mathrm{MSE}(f)=\mathbb{E}[(f-y)^{2}]$. We introduce $g$ by writing $f-y=(f-g)+(g-y)$:
@@
We rewrite the term $2\mathbb{E}[(f-g)(g-y)]$. Since $g-y=(f-y)-(f-g)$, we have:
@@
Before we move on to our analysis of learners in a graph, we extract one more simple but very useful consequence of the decomposition in Lemma 3.3. If a model $f$ is self-orthogonal as well as being multiaccurate with respect to $g$, then not only does $f$ have weakly lower MSE compared to $g$, but the extent to which the actual predictions made by $f$ differ from those made by $g$ can be bounded by the improvement in MSE that $f$ obtains over $g$. Thus two such predictors with similar MSE must actually make very similar predictions.
@@ 3.3 Performance Guarantee under a Random Allocation Model
Let the total set of features be $[d]=\{1,2,\ldots,d\}$. We now assume that for each agent $A_{i}$ in the DAG, its feature set $S_{i}$ is generated randomly. Specifically, for each feature $k\in[d]$, we have $\mathbb{P}[k\in S_{i}]=p$ independently for all $i$ and $k$.
@@
Let’s fix a single subsequence of length $M$ along $P$. Let’s also fix a single feature $k\in[d]$. The probability that a single agent does not observe feature $k$ is $1-p$. Since feature allocations are independent across agents, the probability that none of the $M$ agents in the subsequence observe feature $k$ is $(1-p)^{M}$.
@@ 4 Extension to General Function Classes
In this section, we extend our analysis beyond linear predictors. We show that the core results—monotonic error reduction and the ability to compete with a strong benchmark—can be preserved under this more flexible model. This allows the framework to be applied to a wider range of learning algorithms, such as decision trees or neural networks, operating on local feature subsets. Much of the analysis proceeds identically to Section 3, and where this is the case, we provide explicit references back to the linear analogues and defer full proofs to Appendix A.
@@ 5.1.1 Analysis of the Predictor Structure
Inductive Step: Assume the claims hold for pass $p-1$. Thus, the predictor at the end of pass $p-1$, $\hat{y}_{p-1,k}$, is a linear combination of variables in $S_{p-1,k}=\{z_{k},\ldots,z_{k-(p-1)}\}$.
Now consider pass $p$. The initial prediction for this pass is $\hat{y}_{p,0}=\hat{y}_{p-1,k}$. For an agent observing feature $x_{i}$ with $i\leq k-p$, the local feature is $x_{i}=z_{i}-z_{i-1}$. The indices $i$ and $i-1$ are both smaller than $k-p+1$. Therefore, $x_{i}$ is uncorrelated with every $z_{j}\in S_{p-1,k}$. The incoming predictor $\hat{y}_{p,i-1}$ (which equals $\hat{y}_{p-1,k}$ for $i=1$) has support $S_{p-1,k}$, so it is also uncorrelated with $x_{i}$. The update is trivial: $\hat{y}_{p,i}=\hat{y}_{p,i-1}$. This proves (1) for pass $p$.
@@
Thus, the support of the new predictor is $S_{p,k-p+1}=S_{p-1,k}\cup\{z_{k-p+1},z_{k-p}\}=\{z_{k},\ldots,z_{k-p}\}$. Subsequent updates within pass $p$ (for agents observing features $x_{i}$ with $i>k-p+1$) will only involve features $x_{i}$ constructed from latent variables already in this new support set. Therefore, no new variables are introduced, and the support set remains unchanged for the rest of the pass. At the end of the pass, the support is $S_{p,k}=\{z_{k},\ldots,z_{k-p}\}$. This proves (3) for pass $p$. ∎
@@ 5.1.2 Lower Bound on Mean Squared Error
Let $m=k-j+1$. The inverse of this $m\times m$ matrix has entries $(C^{-1})_{ab}=\frac{1}{m+1}\min(a,b)(m+1-\max(a,b))$. We need the $(1,1)$ entry:
@@
The minimum possible MSE for any such predictor is given by Lemma 5.6, with $j=k-p+1$. Substituting this into the formula gives:
@@ 5.2 Learning Requires Depth in Arbitrary DAGs
Inductive Step: Assume the claim holds for all nodes with depth up to $p-1$. Let $v$ be a node with $\text{depth}(v)=p$. All parents $u\in\text{Pa}(v)$ must have $\text{depth}(u)\leq p-1$. By the inductive hypothesis, for any parent $u$, $\text{Info}(\hat{y}_{u})\subseteq\{z_{k},\ldots,z_{k-(p-1)}\}$.
@@ Acknowledgments
We used Gemini 2.5 Pro within the Windsurf Cascade environment as an aid in writing code and proving lemmas, with detailed instructions from the authors. AI tools were not used for exposition or discussion of related work.
@@ B.2 Complexity Measures and Concentration
^$\hat{\mathcal{R}}_{Z}(\mathcal{G})=\mathbb{E}_{\sigma_{1},\ldots,\sigma_{m}}\left[\sup_{g\in\mathcal{G}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}g(z_{i})\right],$
@@ C.2 Standard Generalization Bound for Linear Predictors
Step 1: Symmetrization and High-Probability Bound. Let =$\mathcal{L}=\{(x,y)\mapsto(f(x)-y)^{2}\mid f\in\mathcal{H}_{\Lambda_{1}}\}$ be the loss class. The proof first relates the expected generalization error to the Rademacher complexity of this loss class, and then uses a concentration inequality to obtain a high-probability bound.
@@
Step 2: Contraction Principle. We bound the complexity of the loss class $\mathcal{R}_{m}(\mathcal{L})$ using the complexity of the hypothesis class $\mathcal{H}_{\Lambda_{1}}$. The loss function can be seen as a composition $\ell(f(x),y)=\phi(f(x)-y)$, where $\phi(z)=z^{2}$. The function $\phi$ is Lipschitz on the interval $[-M_{L},M_{L}]$, with a Lipschitz constant $L_{\phi}=\sup_{z\in[-M_{L},M_{L}]}|\phi^{\prime}(z)|=\sup_{z\in[-M_{L},M_{L}]}|2z|=2M_{L}$.
@@
Let $v=\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}x_{i}$. For each coordinate $j\in\{1,\dots,d\}$, $v_{j}=\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}x_{i,j}$ is an average of independent random variables with mean 0, bounded in $[-X_{\max},X_{\max}]$. By Hoeffding’s inequality, $\mathbb{P}(|v_{j}|\geq t)\leq 2\exp(-mt^{2}/(2X_{\max}^{2}))$. Using a union bound over all $d$ dimensions and a standard bounding argument for the expectation, we get:

### v2

@@ 1.1 Our Results and Techniques
    This is what Gopalan et al. (2022) call “degree 2 calibration”.
We use the fact that a predictor $f$ that is multiaccurate with respect to a collection of functions $G$ and also self-orthogonal has mean squared error that is at most that of the optimal function in $G$ — this is a consequence of the simple MSE decomposition that we state in Lemma 3.3:
@@
and is closely related to a theorem from Gopalan et al. (2022) showing that “degree 2 multicalibration” suffices for error optimality within a class $G$ in squared error. Our requirements are slightly weaker [^1] In fact, their proof (Propostion A.1 of Gopalan et al. (2022)) also only uses these two conditions. We call out the weaker statement here explicitely because these weaker properties — but not full degree 2 multicalibration — follow “for free” from linear regression, which is important for our application. — multiaccuracy is what Gopalan et al. (2022) call degree 1 multicalibration, and so we need only “degree 1 multicalibration” and “degree 2 calibration”, echoing recent work on relaxing multicalibration requirements to “calibrated multiaccuracy” requirements (Gopalan et al., 2023; Casacuberta et al., 2025).
@@
Now recall that one of the inputs to agent $i$’s model $f_{i}$ is the prediction $f_{j}$ of any agent $j$ that is a parent of $i$ in the graph. Hence, it follows from $f_{i}$’s multiaccuracy guarantee that $\mathbb{E}[f_{j}(f_{i}-y)]=0$. Lemma 3.8 shows that this condition implies that:
@@ 1.2 Additional Related Work
Banerjee (1992) and Bikhchandani et al. (1992) study simple models in which Bayesian agents act sequentially, each taking an action that maximizes their utility in expectation over their posterior beliefs of a single, static state. In these models, agents observe the actions of all previous agents, and the main phenomenon of interest is “herding” — i.e. when an infinite sequence of agents takes the wrong action, as the cumulative decisions of their predecessors overwhelm the strength of their own signal. An interesting feature of our model is that herding/information cascades cannot occur indefinitely — this is a consequence of our information aggregation theorems, and a qualitative distinction between our work and models that study binary action games. Acemoglu et al. (2011) extend the earlier herding results to general networks in which agents do not observe the actions of all predecessors, but only those adjacent to them in the network. They study conditions (expansion properties) under which information aggregation occurs in the infinite limit. Mossel et al. (2015, 2020) study a variant of this style of model in which agents act repeatedly rather than sequentially, and study information aggregation in equilibrium. Jadbabaie et al. (2012); Shahrampour and Jadbabaie (2013) study a related model in which agents make Bayesian updates with respect to their own signals, but then broadcast their posterior and simply average their posterior beliefs with those of their neighbors, in the style of a DeGroot (1974) model. They similarly give conditions sufficient for information aggregation.
@@
Technically, our analysis is most closely related to a recent line of work that uses conditional calibration conditions for ensembling and information aggregation Roth et al. (2023); Globus-Harris et al. (2023, 2024); Collina et al. (2025a, b). Two recent works of Collina et al. (2025b, a) in particular show how to recover the same quantitative results in the style of Aumann’s agreement theorem (Aumann, 1976; Aaronson, 2005) and its information aggregation properties Frongillo et al. (2023) without assuming full Bayesian rationality, but instead assuming only simple calibration conditions that can be tractably enforced even in sequential adversarial settings (Noarov et al., 2023; Arunachaleswaran et al., 2025). Our work is distinguished from this literature in two ways: first, this literature considers agents who engage in an extended interaction in which each agent participates multiple times, solving many learning problems across this interaction. In contrast, in our model, each agent learns only once, and then fixes their model: information aggregation needs to occur as predictions percolate through a network of agents, no pair of which might observe the same set of features. Second, this literature assumes stronger conditional calibration conditions (“multi-calibration” (Hébert-Johnson et al., 2018)) that require learning more complicated non-linear models, even to compete with linear benchmark classes. In contrast, our analysis requires only much weaker orthogonality conditions (also known as “multi-accuracy” (Hébert-Johnson et al., 2018; Kim et al., 2019)) which not only are satisfiable by linear models but are already satisfied by least squares regression, and so need no further algorithmic intervention. Our model could be used to simulate the repeated interaction studied in Collina et al. (2025a, b) by constructing a path graph in which each agent was repeatedly represented by alternating nodes — for squared error minimization over additively separable classes in the batch setting, our results would recover some of those of Collina et al. (2025a) with substantially simpler learning algorithms (note however that Collina et al. (2025a) also studies more complex settings with abstract action spaces and utility functions, the online adversarial setting, and a more general weak learning condition).
@@ 2 Preliminaries
Let $d]=\{1,2,\ldots,d\}$ be the set of indices for $d$ total features. Each agent $A_{i}\in\mathcal{A}$ is associated with a specific subset of these features, $S_{i}\subseteq[d]$. We assume data points $(x,y)$ are drawn from an underlying distribution $\mathcal{D}$, where $x\in\mathbb{R}^{d}$ is a feature vector and $y\in\mathbb{R}$ is the target label. For any agent $A_{i}$, its local view of the features is $x_{S_{i}}$, which is the sub-vector of $x$ corresponding to the features indexed by $S_{i}$.
@@
This means that the predictor’s errors f(x)yf(x)-y have low correlation with every function $g$ in the collection $\mathcal{G}$. If $\epsilon=0$ we simply say that $f$ is multiaccurate with respect to $\mathcal{G}$.
@@
Self orthogonality is what Gopalan et al. (2023) define as “degree 2 calibration”.
@@ 3 Linear Learners
Throughout our analysis, we work in the distributional regime (i.e. in the population or infinite-sample limit), assuming that agents can directly optimize over the true distribution $\mathcal{D}$. This allows us to isolate the structural aspects of the equilibrium and learning dynamics. Specifically, under this assumption, linear regression produces predictors that satisfy exact multiaccuracy and exact self-orthogonality, the two key properties that drive our analysis. The finite-sample and out-of-sample generalization guarantees corresponding to these results are established separately in Appendices BC.
@@ 3.1 Structural Properties of Squared Error Regression
Next we observe a useful identity that we will make use of several times: for any pair of predictors $f$ and $g$, we can express the MSE of $f$ in terms of:
@@
This decomposition does not depend on linearity.
@@
We start by expanding $\mathrm{MSE}(f)=\mathbb{E}[(f-y)^{2}]$. We introduce $g$ by writing fy=(fg)+(gy)f-y=(f-g)+(g-y):
@@
We rewrite the term $2\mathbb{E}[(f-g)(g-y)]$. Since gy=(fy)(fg)g-y=(f-y)-(f-g), we have:
@@
Before we move on to our analysis of learners in a graph, we extract one more simple but very useful consequence of the decomposition in Lemma 3.3. If a model $f$ is self-orthogonal as well as being multiaccurate with respect to $g$, then not only does $f$ have weakly lower MSE compared to $g$, but the extent to which the actual predictions made by $f$ differ from those made by $g$ can be bounded by the improvement in MSE that $f$ obtains over $g$. Thus two such predictors with similar MSE must actually make very similar predictions. This next lemma can be viewed as a“pythagorean theorem” for orthogonal projections and is a corollary of other similar pythagorean theorems — e.g. Gopalan et al. (2022) gives one for “degree 2 multicalibration” — but we give the (very simple) proof here for completeness.
@@ 3.3 Performance Guarantee under a Random Allocation Model
Let the total set of features be $d]=\{1,2,\ldots,d\}$. We now assume that for each agent $A_{i}$ in the DAG, its feature set $S_{i}$ is generated randomly. Specifically, for each feature $k\in[d]$, we have $\mathbb{P}[k\in S_{i}]=p$ independently for all $i$ and $k$.
@@
Let’s fix a single subsequence of length $M$ along $P$. Let’s also fix a single feature $k\in[d]$. The probability that a single agent does not observe feature $k$ is 1p1-p. Since feature allocations are independent across agents, the probability that none of the $M$ agents in the subsequence observe feature $k$ is $(1-p)^{M}$.
@@ 4 Extension to General Function Classes
In this section, we extend our analysis beyond linear predictors. We show that the core results—monotonic error reduction and the ability to compete with a strong benchmark—can be preserved under this more flexible model. This allows the framework to be applied to a wider range of learning algorithms, such as decision trees or neural networks, operating on local feature subsets. Much of the analysis proceeds identically to Section 3, and where this is the case, we provide explicit references back to the linear analogues and defer full proofs to Appendix A. As in Section 3, we work in the distributional regime; see Appendix D for finite-sample guarantees.
@@ 5.1.1 Analysis of the Predictor Structure
Inductive Step: Assume the claims hold for pass p1p-1. Thus, the predictor at the end of pass p1p-1, $\hat{y}_{p-1,k}$, is a linear combination of variables in $S_{p-1,k}=\{z_{k},\ldots,z_{k-(p-1)}\}$.
Now consider pass $p$. The initial prediction for this pass is $\hat{y}_{p,0}=\hat{y}_{p-1,k}$. For an agent observing feature $x_{i}$ with $i\leq k-p$, the local feature is $x_{i}=z_{i}-z_{i-1}$. The indices $i$ and i1i-1 are both smaller than kp+1k-p+1. Therefore, $x_{i}$ is uncorrelated with every $z_{j}\in S_{p-1,k}$. The incoming predictor $\hat{y}_{p,i-1}$ (which equals $\hat{y}_{p-1,k}$ for $i=1$) has support $S_{p-1,k}$, so it is also uncorrelated with $x_{i}$. The update is trivial: $\hat{y}_{p,i}=\hat{y}_{p,i-1}$. This proves (1) for pass $p$.
@@
Thus, the support of the new predictor is $S_{p,k-p+1}=S_{p-1,k}\cup\{z_{k-p+1},z_{k-p}\}=\{z_{k},\ldots,z_{k-p}\}$. Subsequent updates within pass $p$ (for agents observing features $x_{i}$ with i>kp+1i>k-p+1) will only involve features $x_{i}$ constructed from latent variables already in this new support set. Therefore, no new variables are introduced, and the support set remains unchanged for the rest of the pass. At the end of the pass, the support is $S_{p,k}=\{z_{k},\ldots,z_{k-p}\}$. This proves (3) for pass $p$. ∎
@@ 5.1.2 Lower Bound on Mean Squared Error
Let m=kj+1m=k-j+1. The inverse of this $m\times m$ matrix has entries $(C^{-1})_{ab}=\frac{1}{m+1}\min(a,b)(m+1-\max(a,b))$. We need the $(1,1)$ entry:
@@
The minimum possible MSE for any such predictor is given by Lemma 5.6, with j=kp+1j=k-p+1. Substituting this into the formula gives:
@@ 5.2 Learning Requires Depth in Arbitrary DAGs
Inductive Step: Assume the claim holds for all nodes with depth up to p1p-1. Let $v$ be a node with $\text{depth}(v)=p$. All parents $u\in\text{Pa}(v)$ must have $\text{depth}(u)\leq p-1$. By the inductive hypothesis, for any parent $u$, $\text{Info}(\hat{y}_{u})\subseteq\{z_{k},\ldots,z_{k-(p-1)}\}$.
@@ Acknowledgments
The authors gratefully acknowledge support from NSF grant CCF-2217062, the NSF Graduate Research Fellowship (grant DGE-2139899), and a grant from the Simons Foundation. We used Gemini 2.5 Pro within the Windsurf Cascade environment as an aid in writing code and proving lemmas, with detailed instructions from the authors. AI tools were not used for exposition or discussion of related work.
@@ B.2 Complexity Measures and Concentration
$\hat{\mathcal{R}}_{Z}(\mathcal{G})=\mathbb{E}_{\sigma_{1},\ldots,\sigma_{m}}\left[\sup_{g\in\mathcal{G}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}g(z_{i})\right],$
@@ C.2 Standard Generalization Bound for Linear Predictors
Step 1: Symmetrization and High-Probability Bound. Let $\mathcal{L}=\{(x,y)\mapsto(f(x)-y)^{2}\mid f\in\mathcal{H}_{\Lambda_{1}}\}$ be the loss class. The proof first relates the expected generalization error to the Rademacher complexity of this loss class, and then uses a concentration inequality to obtain a high-probability bound.
@@
Step 2: Contraction Principle. We bound the complexity of the loss class $\mathcal{R}_{m}(\mathcal{L})$ using the complexity of the hypothesis class $\mathcal{H}_{\Lambda_{1}}$. The loss function can be seen as a composition $\ell(f(x),y)=\phi(f(x)-y)$, where $\phi(z)=z^{2}$. The function $\phi$ is Lipschitz on the interval $M_{L},M_{L}]$, with a Lipschitz constant $L_{\phi}=\sup_{z\in[-M_{L},M_{L}]}|\phi^{\prime}(z)|=\sup_{z\in[-M_{L},M_{L}]}|2z|=2M_{L}$.
@@
Let $v=\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}x_{i}$. For each coordinate $j\in\{1,\dots,d\}$, $v_{j}=\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}x_{i,j}$ is an average of independent random variables with mean 0, bounded in $X_{\max},X_{\max}]$. By Hoeffding’s inequality, $\mathbb{P}(|v_{j}|\geq t)\leq 2\exp(-mt^{2}/(2X_{\max}^{2}))$. Using a union bound over all $d$ dimensions and a standard bounding argument for the expectation, we get: